# On RoCK blocks of double covers of symmetric and alternating groups and the refined Broué conjecture

### v1

@@ 5 On spin blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$
$\mathbb{F}_{p}N^{\rho,d}f\cong B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}).$
Let $P$ be a Sylow $p$-subgroup of $\widetilde{S}_{p}$. The block algebra $\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1}$ has a cyclic defect group $P$ and Brauer correspondent $\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0}$. By Lemma 6.4, there is a splendid Rickard complex $X$ of $((B^{\varnothing,1}_{\mathbb{F}_{p}})_{\bar{0}},(b^{\varnothing,1}_{\mathbb{F}_{p}})_{\bar{0}})$-bimodule that extends to a complex $\tilde{X}$ of $(B^{\varnothing,1}_{\mathbb{F}_{p}},b^{\varnothing,1}_{\mathbb{F}_{p}})_{C_{2}}$-modules. Let $H$ be the subgroup of $G$ generated by $\widetilde{S}_{U}$ and all $\widetilde{A}_{V_{i}}$’s. Now (by Propositions 4.2, 4.12 and Remark 2.3) $X^{\otimes d}$ induces a splendid Rickard equivalence between
$\mathbb{F}_{p}(H\cap\widetilde{S}_{V})e_{\varnothing,1}^{\otimes d}\cong\mathbb{F}_{p}\widetilde{A}_{V_{1}}e_{\varnothing,1}\otimes_{\mathbb{F}_{p}}\cdots\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}\widetilde{A}_{V_{d}}e_{\varnothing,1}\cong(B^{\varnothing,1}_{\bar{0}})^{\otimes d}$
@@
$\mathbb{F}_{p}(N_{G}(D)\cap H\cap\widetilde{S}_{V})e_{\varnothing,0}\cong\mathbb{F}_{p}N_{\widetilde{A}_{V_{1}}}(D_{1})e_{\varnothing,0}\otimes_{\mathbb{F}_{p}}\cdots\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}N_{\widetilde{A}_{V_{d}}}(D_{d})e_{\varnothing,0}\cong(b^{\varnothing,1}_{\bar{0}})^{\otimes d}.$
We now adopt the identifications (6.3). By Proposition 2.12, $\tilde{X}^{\otimes d}$ extends to a complex $Y_{0}$ of modules for the group algebra of
$\langle(H\cap\widetilde{S}_{V})\times(N_{G}(D)\cap H\cap\widetilde{S}_{V}),(u_{i},u_{i}),(T_{w},T_{w})\rangle\leq(N^{\rho,d}\cap\widetilde{S}_{V})\times(N_{G}(D)\cap\widetilde{S}_{V}),$
where $i$ ranges over $\{1,\cdots,d\}$ with $u_{i}\in N_{\widetilde{S}_{V_{i}}}(D_{i})\backslash N_{\widetilde{A}_{V_{i}}}(D_{i})$, and $w$ ranges over $S_{d}$ with $T_{w}$ defined as in 5.12. By Proposition 4.7 (iii), $Y_{0}$ is a splendid complex. We then induce $Y_{0}$ up to
$(N^{\rho,d}\cap\widetilde{S}_{V})\times(N_{G}(D)\cap\widetilde{S}_{V})$
@@
$\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}_{V})e_{\varnothing,1}^{\otimes d}\cong B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}$
@@
$\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}_{V})e_{\varnothing,0}^{\otimes d}\cong b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}.$
By Proposition 4.7 (i), $Y$ is a splendid complex.
@@
for $1\leq i\leq d$ and $1\leq j\leq d-1$, where $g_{i}$ is the generator of the $i$-th $C_{2}$ and $s_{j}$ is the appropriate elementary transposition in $S_{d}$. Let $\mathcal{K}$ be the kernel of this homomorphism. Since $B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}$ is a $C_{2}\wr S_{d}$-graded algebra, we use the notation $(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})_{g}$ to denote the $g$-component of $B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}$ for any $g\in C_{2}\wr S_{d}$. Then we have
$(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})_{\bar{0}}=\bigoplus_{g\in\mathcal{K}}(B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}})_{g}.$
@@
$\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}_{V})e_{\varnothing,1}^{\otimes d}\cong B^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}$
@@
$\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}_{V})e_{\varnothing,0}^{\otimes d}\cong b^{\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}_{d}^{\mathbb{F}_{p}}.$
It is straightforward to check that the $(B^{\rho,0}_{\mathbb{F}_{p}},B^{\rho,0}_{\mathbb{F}_{p}})$-bisupermodule $B^{\rho,0}_{\mathbb{F}_{p}}$ induces a Morita self-superequivalence of $B^{\rho,0}_{\mathbb{F}_{p}}$; this bimodule is splendid because it is isomorphic to a direct summand of $\mathbb{F}_{p}\widetilde{S}_{r}\cong{\rm Ind}_{\Delta\widetilde{S}_{r}}^{\widetilde{S}_{r}\times\widetilde{S}_{r}}(\mathbb{F}_{p})$. Hence by Proposition 4.2, the complex $B^{\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y$ of $\mathbb{Z}/2\mathbb{Z}$-graded $(B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}_{V})e_{\rho,1}^{\otimes d})$-$(B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}_{V})e_{\rho,0}^{\otimes d})$-bimodule induces a $\mathbb{Z}/2\mathbb{Z}$-graded Rickard equivalence between
$\mathbb{F}_{p}N^{\rho,d}f\cong B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}_{V})e_{\varnothing,1}^{\otimes d}$
@@
$\mathbb{F}_{p}N_{G}(D)e_{\rho,d}\cong B^{\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}_{V})e_{\varnothing,0}^{\otimes d}.$
By Proposition 4.12, $B^{\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y$ is a splendid complex of $(\mathbb{F}_{p}N^{\rho,d}f,\mathbb{F}_{p}N_{G}(D)e_{\rho,d})$-bimodules.
Now by Theorem 4.4, $(B^{\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y)_{\bar{0}}$ induces a Rickard equivalence between $\mathbb{F}_{p}N^{\rho,d}_{\bar{0}}f$ and $\mathbb{F}_{p}N_{G}(D)_{\bar{0}}e_{\rho,d}$. Since $(B^{\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y)_{\bar{0}}$ is isomorphic to a direct summand of the complex
${\rm Res}_{N^{\rho,d}_{\bar{0}}\times N_{G}(D)_{\bar{0}}}^{N^{\rho,d}\times N_{G}(D)}(B^{\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y),$
@@
7.1.
Throughout this section we set $d$ to be an integer with $1\leq d<p$ and $\rho$ a $d$-Rouquier $\bar{p}$-core. As in Section 5, we continue to assume that $K$ (and hence $\mathcal{O}$) contains a primitive $(2n!)$-th root of unity in this section. We also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$); this condition is automatic if $n\geq 4$. Let $R\in\{\mathcal{O},k\}$ and $R_{p}\in\{\mathbb{Z}_{p},\mathbb{F}_{p}\}$. We will adopt the notation in 5.13, in particular, $r=|\rho|$, $n=r+dp$,
$G=\widetilde{S}_{n},\quad L=\widetilde{S}_{U,V_{1},\dots,V_{d}},\quad N^{\rho,d}=N_{G}(\widetilde{S}_{V_{1},\cdots,V_{d}}),\quad$
@@
is chosen as in Lemma 5.15. We identify $RLf$ with $B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R})^{\otimes d}$ and $RN^{\rho,d}f$ with $B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})$ via the isomorphisms (5.8). Let $\ell:=(p-1)/2$.
7.2.
The bisupermodule which induces a Morita superequivalence. We set $M^{R}$ to be the $(B^{\varnothing,1}_{R},B^{\varnothing,1}_{R})$-bisupermodule
$M^{R}:=\Omega_{B^{\varnothing,1}_{R}\otimes_{R}(B^{\varnothing,1}_{R})^{\rm sop}}^{\ell}(B^{\varnothing,1}_{R}).$
@@
Then we have the $((B^{\varnothing,1}_{R})^{\otimes d},(B^{\varnothing,1}_{R})^{\otimes d})$-bisupermodule $(M^{R})^{\boxtimes d}$. For $1\leq i\leq d$, we identify $(B^{\varnothing,1}_{R})^{(i)}:=R\widetilde{S}_{V_{i}}e_{\varnothing,1}^{(i)}$ with $B^{\varnothing,1}_{R}=R\widetilde{S}_{p}e_{\varnothing,1}$ via (5.5), as in 5.12. Denote by $(M^{R})^{(i)}$ the $i$-th factor in $(M^{R})^{\boxtimes d}$. That is, $(M^{R})^{(i)}$ is a $((B^{\varnothing,1}_{R})^{(i)},(B^{\varnothing,1}_{R})^{(i)})$-bisupermodule that we identify with $M^{R}$. Denote by $(M^{R})_{S_{d}}^{\boxtimes d}$ the $(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R},B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})_{S_{d}}$-supermodule defined as in (2.5).
Recalling the identification $RLf\cong B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R})^{\otimes d}$, define the $(RLf,RLf)$-bisupermodule
$M_{L}^{R}:=B^{\rho,0}_{R}\boxtimes_{R}(M^{R})^{\boxtimes d}.$
@@
We have also identified $RN^{\rho,d}f$ with $B^{\rho,0}_{R}\otimes_{R}(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})$. Recalling (2.6), define the $(RN^{\rho,d}f,RN^{\rho,d}f)$-bisupermodule
$M_{N^{\rho,d}}^{R}:=B^{\rho,0}_{R}\boxtimes_{R}(M^{R}\wr_{s}\mathcal{T}_{d}^{R})=B^{\rho,0}_{R}\boxtimes_{R}\Big({\rm Ind}_{(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R},B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})_{S_{d}}}^{(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})\otimes_{R}(B^{\varnothing,1}_{R}\wr_{s}\mathcal{T}_{d}^{R})^{\rm sop}}(M^{R})^{\boxtimes d}_{S_{d}}\Big).$
@@
By [22, Lemma 7.2 (ii)], the $(RN^{\rho,d}f,RN^{\rho,d}f)$-bimodule $M_{N^{\rho,d}}^{R}$ is module-indecomposable with vertex $\Delta D:=\{(u,u)\mid u\in D\}$, and by [22, Lemma 4.8 (ii)], we have $N_{G\times N^{\rho,d}}(\Delta D)\leq N^{\rho,d}\times N^{\rho,d}$. Hence we can now define $X^{R}$ to be the super Green correspondent of $M_{N^{\rho,d}}^{R}$ in $G\times N^{\rho,d}$; see Theorem 3.11 for the definition of the super Green correspondence. By [22, Lemma 7.5], $X^{R}$ is an $(RGb,RN^{\rho,d}f)$-bisupermodule.
Theorem 7.3 ([22, Theorem 8.13]).
@@
The bisupermodules $X^{R}$ and $(X^{R})^{*}$ induce a Morita superequivalence between $RGb$ and $RN^{\rho,d}f$. $X^{R}_{\bar{0}}$ and $(X^{R}_{\bar{0}})^{*}$ induce a Morita superequivalence between $RG_{\bar{0}}b$ and $RN_{\bar{0}}^{\rho,d}f$. Both the bimodules $X^{R}$ and $X^{R}_{\bar{0}}$ have endopermutation sources.

### v5

@@ 6 On spin blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$
$\mathbb{F}_{p}N^{\rho,d}f\cong B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}(B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}).$
Let $P$ be a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{p}$. The block algebra $\mathbb{F}_{p}\widetilde{A}_{p}e_{\varnothing,1}$ has a cyclic defect group $P$ and Brauer correspondent $\mathbb{F}_{p}N_{\widetilde{A}_{p}}(P)e_{\varnothing,0}$. By Lemma 7.4, there is a splendid Rickard complex $X$ of $((B^{\eta,\varnothing,1}_{\mathbb{F}_{p}})_{\bar{0}},(b^{\eta,\varnothing,1}_{\mathbb{F}_{p}})_{\bar{0}})$-bimodule that extends to a complex $\tilde{X}$ of $(B^{\eta,\varnothing,1}_{\mathbb{F}_{p}},b^{\eta,\varnothing,1}_{\mathbb{F}_{p}})_{C_{2}}$-modules. Let $H$ be the subgroup of $G$ generated by $\widetilde{S}^{\eta}_{U}$ and all $\widetilde{A}_{V_{i}}$’s. Now (by Propositions 5.2, 5.14 and Remark 2.3) $X^{\otimes d}$ induces a splendid Rickard equivalence between
$\mathbb{F}_{p}(H\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,1}^{\otimes d}\cong\mathbb{F}_{p}\widetilde{A}_{V_{1}}e_{\varnothing,1}\otimes_{\mathbb{F}_{p}}\cdots\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}\widetilde{A}_{V_{d}}e_{\varnothing,1}\cong(B^{\eta,\varnothing,1}_{\bar{0}})^{\otimes d}$
@@
$\mathbb{F}_{p}(N_{G}(D)\cap H\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,0}\cong\mathbb{F}_{p}N_{\widetilde{A}_{V_{1}}}(D_{1})e_{\varnothing,0}\otimes_{\mathbb{F}_{p}}\cdots\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}N_{\widetilde{A}_{V_{d}}}(D_{d})e_{\varnothing,0}\cong(b^{\eta,\varnothing,1}_{\bar{0}})^{\otimes d}.$
We now adopt the identifications (7.4). By Proposition 3.8, $X^{\otimes d}$ extends to a complex $Y_{0}$ of modules for the group algebra of
$\langle(H\cap\widetilde{S}^{\eta}_{V})\times(N_{G}(D)\cap H\cap\widetilde{S}^{\eta}_{V}),(u_{i},u_{i}),(T_{w},T_{w})\rangle\leq(N^{\rho,d}\cap\widetilde{S}^{\eta}_{V})\times(N_{G}(D)\cap\widetilde{S}^{\eta}_{V}),$
where $i$ ranges over $\{1,\cdots,d\}$ with $u_{i}\in N_{\widetilde{S}^{\eta}_{V_{i}}}(D_{i})\backslash N_{\widetilde{A}_{V_{i}}}(D_{i})$, and $w$ ranges over $S_{d}$ with $T_{w}$ chosen as in Lemma 6.12. We then induce $Y_{0}$ up to
$(N^{\rho,d}\cap\widetilde{S}^{\eta}_{V})\times(N_{G}(D)\cap\widetilde{S}^{\eta}_{V})$
@@
$\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,1}^{\otimes d}\cong B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}$
@@
$\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,0}^{\otimes d}\cong b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}.$
By Proposition 5.6 (iii), $Y$ is a splendid complex.
@@
for $1\leq i\leq d$ and $1\leq j\leq d-1$, where $g_{i}$ is the generator of the $i$-th $C_{2}$ and $s_{j}$ is the appropriate elementary transposition in $S_{d}$. Let $\mathcal{K}$ be the kernel of this homomorphism. Since $B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}$ is a $C_{2}\wr S_{d}$-graded algebra, we use the notation $(B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}})_{g}$ to denote the $g$-component of $B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}$ for any $g\in C_{2}\wr S_{d}$. Then we have
$(B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}})_{\bar{0}}=\bigoplus_{g\in\mathcal{K}}(B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}})_{g}.$
@@
$\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,1}^{\otimes d}\cong B^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}$
@@
$\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,0}^{\otimes d}\cong b^{\eta,\varnothing,1}_{\mathbb{F}_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,\mathbb{F}_{p}}.$
It is straightforward to check that the $(B^{\eta,\rho,0}_{\mathbb{F}_{p}},B^{\eta,\rho,0}_{\mathbb{F}_{p}})$-bisupermodule $B^{\eta,\rho,0}_{\mathbb{F}_{p}}$ induces a Morita self-superequivalence of $B^{\eta,\rho,0}_{\mathbb{F}_{p}}$; this bimodule is splendid because it is isomorphic to a direct summand of $\mathbb{F}_{p}\widetilde{S}^{\eta}_{r}\cong{\rm Ind}_{\Delta\widetilde{S}^{\eta}_{r}}^{\widetilde{S}^{\eta}_{r}\times\widetilde{S}^{\eta}_{r}}(\mathbb{F}_{p})$. Hence by Proposition 5.2, the complex $B^{\eta,\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y$ of $\mathbb{Z}/2\mathbb{Z}$-graded $(B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,1}^{\otimes d})$-$(B^{\eta,\varnothing,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}^{\eta}_{V})e_{\rho,0}^{\otimes d})$-bimodule induces a $\mathbb{Z}/2\mathbb{Z}$-graded Rickard equivalence between
$\mathbb{F}_{p}N^{\rho,d}f\cong B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N^{\rho,d}\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,1}^{\otimes d}$
@@
$\mathbb{F}_{p}N_{G}(D)e_{\rho,d}\cong B^{\eta,\rho,0}_{\mathbb{F}_{p}}\otimes_{\mathbb{F}_{p}}\mathbb{F}_{p}(N_{G}(D)\cap\widetilde{S}^{\eta}_{V})e_{\varnothing,0}^{\otimes d}.$
By Proposition 5.14, $B^{\eta,\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y$ is a splendid complex of $(\mathbb{F}_{p}N^{\rho,d}f,\mathbb{F}_{p}N_{G}(D)e_{\rho,d})$-bimodules.
Now by Theorem 5.3, $(B^{\eta,\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y)_{\bar{0}}$ induces a Rickard equivalence between $\mathbb{F}_{p}N^{\rho,d}_{\bar{0}}f$ and $\mathbb{F}_{p}N_{G_{\bar{0}}}(D)e_{\rho,d}$. Since $(B^{\eta,\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y)_{\bar{0}}$ is isomorphic to a direct summand of the complex
${\rm Res}_{N^{\rho,d}_{\bar{0}}\times N_{G_{\bar{0}}}(D)}^{N^{\rho,d}\times N_{G}(D)}(B^{\eta,\rho,0}_{\mathbb{F}_{p}}\boxtimes_{\mathbb{F}_{p}}Y),$
@@
8.1.
Throughout this section we set $d$ to be an integer with $1\leq d<p$ and $\rho$ a $d$-Rouquier $\bar{p}$-core. As in Section 6, we continue to assume that $K$ (and hence $\mathcal{O}$) contains a primitive $(2n!)$-th root of unity in this section. We also assume that 1-1 and $2$ have square roots in $K$ (and hence $\mathcal{O}$); this condition is automatic if $n\geq 4$. Let $(R,R_{p})\in\{(\mathcal{O},\mathbb{Z}_{p}),(k,\mathbb{F}_{p})\}$. We may assume that $R$ is finitely generated as $R_{p}$-module. We will adopt the notation in 6.13, in particular, $r=|\rho|$, $n=r+dp$,
$G=\widetilde{S}^{\eta}_{n},\quad L=\widetilde{S}^{\eta}_{U,V_{1},\dots,V_{d}},\quad N^{\rho,d}=N_{G}(\widetilde{S}^{\eta}_{V_{1},\cdots,V_{d}}),\quad$
@@
is chosen as in Lemma 6.15. We identify $RLf$ with $B^{\eta,\rho,0}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R})^{\otimes d}$ and $RN^{\rho,d}f$ with $B^{\eta,\rho,0}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R})$ via the isomorphisms (6.8). Let $\ell:=(p-1)/2$.
8.2.
The bisupermodule which induces a Morita superequivalence. We set $M^{R}$ to be the $(B^{\eta,\varnothing,1}_{R},B^{\eta,\varnothing,1}_{R})$-bisupermodule
$M^{R}:=\Omega_{B^{\eta,\varnothing,1}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R})^{\rm sop}}^{\ell}(B^{\eta,\varnothing,1}_{R}).$
@@
Then we have the $((B^{\eta,\varnothing,1}_{R})^{\otimes d},(B^{\eta,\varnothing,1}_{R})^{\otimes d})$-bisupermodule $(M^{R})^{\boxtimes d}$. For $1\leq i\leq d$, we identify $(B^{\eta,\varnothing,1}_{R})^{(i)}:=R\widetilde{S}^{\eta}_{V_{i}}e_{\varnothing,1}^{(i)}$ with $B^{\eta,\varnothing,1}_{R}=R\widetilde{S}^{\eta}_{p}e_{\varnothing,1}$ via (6.5), as in 6.11. Denote by $(M^{R})^{(i)}$ the $i$-th factor in $(M^{R})^{\boxtimes d}$. That is, $(M^{R})^{(i)}$ is a $((B^{\eta,\varnothing,1}_{R})^{(i)},(B^{\eta,\varnothing,1}_{R})^{(i)})$-bisupermodule that we identify with $M^{R}$. Denote by $(M^{R})_{S_{d}}^{\boxtimes d}$ the $(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R},B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R})_{S_{d}}$-supermodule defined as in (3.2).
Recalling the identification $RLf\cong B^{\eta,\rho,0}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R})^{\otimes d}$, define the $(RLf,RLf)$-bisupermodule
$M_{L}^{R}:=B^{\eta,\rho,0}_{R}\boxtimes_{R}(M^{R})^{\boxtimes d}.$
@@
We have also identified $RN^{\rho,d}f$ with $B^{\eta,\rho,0}_{R}\otimes_{R}(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R})$. Recalling (3.3), define the $(RN^{\rho,d}f,RN^{\rho,d}f)$-bisupermodule
$M_{N^{\rho,d}}^{R}:=B^{\eta,\rho,0}_{R}\boxtimes_{R}(M^{R}\wr_{s}\mathcal{T}^{\eta}_{d,R})=B^{\eta,\rho,0}_{R}\boxtimes_{R}\Big({\rm Ind}_{(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R},B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R})_{S_{d}}}^{(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R})\otimes_{R}(B^{\eta,\varnothing,1}_{R}\wr_{s}\mathcal{T}^{\eta}_{d,R})^{\rm sop}}(M^{R})^{\boxtimes d}_{S_{d}}\Big).$
@@
By [23, Lemma 7.2 (ii)] and Proposition 3.6 (i), the $(RN^{\rho,d}f,RN^{\rho,d}f)$-bimodule $M_{N^{\rho,d}}^{R}$ is module-indecomposable with vertex $\Delta D:=\{(u,u)\mid u\in D\}$, and by [23, Lemma 4.8 (ii)], we have $N_{G\times N^{\rho,d}}(\Delta D)\leq N^{\rho,d}\times N^{\rho,d}$. Hence we can now define $X^{R}$ to be the super Green correspondent of $M_{N^{\rho,d}}^{R}$ in $G\times N^{\rho,d}$; see Theorem 4.11 for the definition of the super Green correspondence. By [23, Lemma 7.5], $X^{R}$ is an $(RGb,RN^{\rho,d}f)$-bisupermodule.
Theorem 8.3 ([23, Theorem 8.13 and Lemma 9.2]).
@@
The bisupermodules $X^{R}$ and $(X^{R})^{*}$ induce a Morita superequivalence between $RGb$ and $RN^{\rho,d}f$. The bimodules $X^{R}_{\bar{0}}$ and $(X^{R}_{\bar{0}})^{*}$ induce a Morita equivalence between $RG_{\bar{0}}b$ and $RN_{\bar{0}}^{\rho,d}f$. Both the bimodules $X^{R}$ and $X^{R}_{\bar{0}}$ have endopermutation sources.